OpenStudy (anonymous):

Show the Curves 2x^2+3y^2=5 and y^2=x^2 are orthogonal. I'm less interested in the answer and more interested in how to do it. Please help. Thank you

OpenStudy (amistre64):

interesting .... im sure it has to do with their derivatives being orthogonal at all points (x,y)

OpenStudy (amistre64):

there is also something i recall that an inner product equal to 0 defines orthogonal functions

OpenStudy (anonymous):

i wrote down in my notes that I first have to find the derivative of 2x^2+3y^2=5? which ends up being: -2x/3y. How is that?

OpenStudy (amistre64):

implicit derivative

OpenStudy (amistre64):

\[D_x[2x^2 + 3y^2 = 5]\] \[D_x[2x^2] + D_x[3y^2] = D_x[5]\] \[2D_x[x^2] + 3D_x[y^2] = 0\] \[4xD_x[x] + 6yD_x[y] = 0\] \[4x + 6y~y' = 0\] \[y' = -\frac{4x}{6y}\]

OpenStudy (anonymous):

how did you go from 4x+6yy'=0 to y' = -4x/6y

OpenStudy (anonymous):

Orthogonal curves have orthogonal/perpendicular tangent lines at every point they intersect. Recall that for two perpendicular lines with slopes \(m_1,m_2\) we have \(m_1m_2=-1\Leftrightarrow m_2=-\frac1{m_1}\Leftrightarrow m_1=-\frac1{m_2}\). Now we just need to show that \(y'_1y'_2=-1\). First, we find \(y'_1\) by implicit differentiation (w.r.t \(x\)):$$\frac{d}{dx}[2x^2+3y_1^2]=\frac{d}{dx}[5]\\4x+6y_1y'_1=0\\6y_1y'_1=-4x\\y'_1=-\frac{4x}{6y_1}=-\frac{2x}{3y_1}$$We do something similar for \(y'_2\):$$\frac{d}{dx}[y_2^2]=\frac{d}{dx}[x^2]\\2y_2y'_2=2x\\y'_2=\frac{2x}{2y_2}=\frac{x}{y_2}$$Now we need to show that for any time they intersect, i.e. \(y_1=y_2\), that \(y_1'y_2'=-1\)... unfortunately:$$y_1'y_2'=-\frac{2x}{3y}\times\frac{x}y=-\frac{2x^2}{3y^2}$$Doesn't seem orthogonal to me.

OpenStudy (anonymous):

Ah I see that makes sense thanks for all the baby steps. And what would be the ordered pair then?

OpenStudy (anonymous):

@Andysebb what ordered pair? where they intersect? Just substitute.$$y^2=x^2\\3y^2=3x^2$$so$$2x^2+3y^2=5\\2x^2+3x^2=5\\5x^2=5\\x^2=1$$so they will intersect at \((\pm1,\pm1)\).

OpenStudy (anonymous):

Notice that \(y^2=x^2\) which means our earlier expression for \(y_1'y_2'\) can be simplified to just \(-\frac23\), which means they're not orthogonal.

OpenStudy (amistre64):

4x+6yy'=0 solve for y' by the most basic of mathematical procedures .. subtract 4x from both sides 6yy' = -4x then divide off the 6y y' = -4x/6y and reduce ....